Use the given frequency distribution to find the (a) class width. (b) class midpoints of the first class. (c) class boundaries of the first class. Phone Calls (per day Class | Frequency, 18 12-1523 16-19 20-2347 24-27 32 38 Select one: O A. (a)3 (b) 9,5 (c) 7.5-11.5 B. (a) 4 (b) 9.5 (c) 7.5-11.5 C.(a) 4 (b) 10.5 (c) 8-11 O D. (a) 3 (b) 10.5 (c) 8-11
Answer #1

a.
Class width = 11-8 = 3

b.
Midpoint of the first class = (8+11)/2 = 9.5

c.

class boundaries of the first class after converting into exclusive class interval is
Class boundaries pf the first = 7.5-11.5

Correct option is (A)